OpenStudy (anonymous):

Evaluate 2^2 − 3 ⋅ 6

OpenStudy (anonymous):

i got 4 - 3 x6= 4-9= -5 which is not one of the choices

OpenStudy (anonymous):

2^2=4 3*6=18 4-18=-12

OpenStudy (anonymous):

what r the choices

OpenStudy (anonymous):

-14, 28, 29

OpenStudy (anonymous):

srry 4-18=-14 lol

OpenStudy (anonymous):

okay, ty :))

OpenStudy (anonymous):

tag me if you need any more help

OpenStudy (anonymous):

okay yeah i do need more help, -2 + 4 x 3^2 - 6

OpenStudy (anonymous):

@muneeb180 idk if i have to tag u in this but tEST